# From Bayesian Asymptotics to General Large-Scale MIMO Capacity

### v2

@@ IV-B Approximate Jeffreys constellation
$\Xi:=\left\{\xi\in\mathbb{R}^{d+1}:\int_{-A}^{A}f_{\xi}(\theta)\,d\theta=1,\;\;f_{\xi}(\theta)\geq 0,\;\theta\in[-A,A]\right\}.$
@@
Solving $P_{3}(\gamma)$ for a sequence of $\gamma\to 0$ yields the solution to $P_{2}$. Note that the gradient and Hessian have the following analytical expression:
$\nabla_{\xi}\!\left[D(f_{\xi}\,\|\,w_{J,\lambda^{*}})+\gamma D(\mathbf{1}\,\|\,f_{\xi})\right]$ $=\int_{-A}^{A}\nabla_{\xi}f_{\xi}(\theta)\left[\ln f_{\xi}(\theta)+1+\lambda^{*}c(\theta)-\tfrac{1}{2}\ln J(\theta)-\frac{\gamma}{2A}\,\frac{1}{f_{\xi}(\theta)}\right]d\theta,$ (56)
$\nabla_{\xi}^{2}\!\left[D(f_{\xi}\,\|\,w_{J,\lambda^{*}})+\gamma D(\mathbf{1}\,\|\,f_{\xi})\right]$ $=\int_{-A}^{A}\nabla_{\xi}f_{\xi}(\theta)\,\nabla_{\xi}f_{\xi}(\theta)^{\top}\left[\frac{1}{f_{\xi}(\theta)}+\frac{\gamma}{2A}\frac{1}{f_{\xi}(\theta)^{2}}\right]d\theta.$ (57)
We can verify that the Hessian is positive definite, which implies that $P_{3}(\gamma)$ is strictly convex. All integrals involves above can be evaluated numerically using a sufficiently fine discretization of [A,A][-A,A] and evaluating $c(\theta)$ and $J(\theta)$ at the grid points. In practice, only a few Newton steps are needed for the convergence to the optimal solution $\xi^{*}$ of $P_{3}(\gamma)$. Given the optimal coefficients $\xi^{*}$, the cdf has a closed form:
@@ IV-C Multi-dimensional constellations
Without any additional structure that may reduce complexity, we must approximate for each $i=1,\ldots,d$, the conditional cdf $F({\theta_{i}}\!\mid\!\theta^{i-1})$ for each $\theta^{i-1}\in\mathcal{A}^{(i-1)}$ where $\mathcal{A}^{(i-1)}$ is the (i1)(i-1)-dimensional constellation previously constructed, whose size is $M^{i-1}$. This leads to a total complexity proportionial to $1+M+\cdots+M^{d-1}=\Theta(M^{d})$. If, however, the model exhibits a Markovian structure of order $D$, then the approximation only needs to be performed for each $\theta_{i-D}^{i-1}$ in a $D$-dimensional grid with size $M^{D}$, yielding a reduced complexity of $\Theta(M^{D})$. In the fully general case, as the dimension $d$ grows, the complexity increases exponentially. Indeed, it is well known that sampling accurately from from high dimension distributions is computationally hard in general.
@@ V-C Achievable rate for continuous output alphabet and approximate log-likelihood
$\boldsymbol{J}(Y;{\boldsymbol{\theta}}\!\mid\!\hat{Y}=\hat{y})=\mathbb{E}\Bigl[\nabla_{\boldsymbol{\theta}}\ln p_{\boldsymbol{\theta}}(Y\!\mid\!\hat{Y}=\hat{y})\,\nabla_{\boldsymbol{\theta}}\ln p_{\boldsymbol{\theta}}(Y\!\mid\!\hat{Y}=\hat{y})^{\top}\,\Big|\,\hat{Y}=\hat{y}\Bigr$]. (66)
@@ VII Conclusion
$C(P)$ $\geq I({\boldsymbol{\theta}};\boldsymbol{y})$ (105)
$=\bar{C}(P)+o(1).$ (106)
@@
$C_{L}(P)-D(w_{J,\lambda^{*}}\|w_{J,\lambda^{*}_{L}})+\lambda_{L}^{*}\,\mathbb{E}_{{\boldsymbol{\theta}}\sim w_{J,\lambda^{*}}}\left[c({\boldsymbol{\theta}})-P\right].$ (108)
@@
$\int_{\Theta}2^{-\lambda^{*}(c({\boldsymbol{\theta}})-P)}\sqrt{\det(\boldsymbol{J}({\boldsymbol{\theta}}))}\log{\sqrt{\det(\boldsymbol{J}({\boldsymbol{\theta}}))}\over\sqrt{\det(\boldsymbol{J}_{L}({\boldsymbol{\theta}}))}}d{\boldsymbol{\theta}}\leq\kappa_{0}\int_{\Theta}\ln{{\det(\boldsymbol{J}({\boldsymbol{\theta}}))}\over{\det(\boldsymbol{J}_{L}({\boldsymbol{\theta}}))}}d{\boldsymbol{\theta}},$ (111)
for some bounded $\kappa_{0}$, since $2^{-\lambda^{*}(c({\boldsymbol{\theta}})-P)}\sqrt{\det(\boldsymbol{J}({\boldsymbol{\theta}}))}$ is bounded for compact $\Theta$ and bounded $P$. Plugging (107), (110), and (111) back in (108), we obtain an achievable rate $C(P)-\kappa_{0}e_{L}$ and complete the proof of the theorem. ∎
@@
$\boldsymbol{J}(Y;{\boldsymbol{\theta}}\!\mid\!\hat{Y}=\ell)=\mathbb{E}\Bigl[\nabla_{\boldsymbol{\theta}}\ln p_{\boldsymbol{\theta}}(Y\!\mid\!\hat{Y}=\ell)\,\nabla_{\boldsymbol{\theta}}\ln p_{\boldsymbol{\theta}}(Y\!\mid\!\hat{Y}=\ell)^{\top}\,\Big|\,\hat{Y}$$\ell\Bigr$]. (120)
@@
$=36\rho^{2}\frac{\eta_{1}\epsilon+2(2+\epsilon)}{\eta_{1}\epsilon}\,L^{-\frac{2\eta_{1}\epsilon}{m(\eta_{1}\epsilon+2(2+\epsilon))}}\left(\frac{\eta_{1}\epsilon\,K_{\epsilon}\kappa_{1}^{\epsilon/(2+\epsilon)}}{72(2+\epsilon)\rho^{2}}\right)^{\frac{2(2+\epsilon)}{\eta_{1}\epsilon+2(2+\epsilon)}}.$ (131)
If $\sup_{{\boldsymbol{\theta}}\in\Theta}P_{{\boldsymbol{\theta}}}(\|Y\|>r)\leq\kappa_{2}e^{-\eta_{2}r^{2}}$ then

### v3

@@ III-E Optical SIMO Poisson channel
Note that the beneficial effect of dithering in multi-antenna systems has been previously observed in the context of coarse quantization and generalized mutual information (GMI) analysis (see, e.g., [11, 22]). In contrast, our formulation makes the impact of dithering explicit through the Fisher information and the resulting Jeffreys factor, thereby providing a simple and quantitative characterization of the associated capacity gain within the Bayesian asymptotic framework.
@@ IV-B Approximate Jeffreys constellation
$\Xi:=\left\{\xi\in\mathbb{R}^{d+1}:f_{\xi}\geq 0\text{ on }[-A,A],\;\int_{-A}^{A}f_{\xi}(\theta)\,d\theta=1\right\}.$
@@
Solving $P_{3}(\gamma)$ for a sequence of $\gamma\to 0$ yields the solution to $P_{2}$. Define the objective function $\mathcal{L}(\xi):=D(f_{\xi}\|w_{J,\lambda^{*}})+\gamma D(\mathbf{1}\|f_{\xi})$, and
$\psi_{\xi}(\theta):=\ln f_{\xi}(\theta)+1+\lambda^{*}c(\theta)-\tfrac{1}{2}\ln J(\theta)-\frac{\gamma}{2A}\frac{1}{f_{\xi}(\theta)}.$ (56)
The gradient and Hessian have the following analytical expression:
$\nabla_{\xi}\mathcal{L}(\xi)$ $=\int_{-A}^{A}\nabla_{\xi}f_{\xi}(\theta)\,\psi_{\xi}(\theta)\,d\theta,$ (57)
$\nabla_{\xi}^{2}\mathcal{L}(\xi)$ $=\int_{-A}^{A}\nabla_{\xi}f_{\xi}(\theta)\nabla_{\xi}f_{\xi}(\theta)^{\top}\left(\frac{1}{f_{\xi}(\theta)}+\frac{\gamma}{2A}\frac{1}{f_{\xi}(\theta)^{2}}\right)d\theta.$ (58)
We can verify that the Hessian is positive definite, which implies that $P_{3}(\gamma)$ is strictly convex. All integrals involved above can be evaluated numerically using a sufficiently fine discretization of [A,A][-A,A] and evaluating $c(\theta)$ and $J(\theta)$ at the grid points. In practice, only a few Newton steps are needed for the convergence to the optimal solution $\xi^{*}$ of $P_{3}(\gamma)$. Given the optimal coefficients $\xi^{*}$, the cdf has a closed form:
@@ IV-C Multi-dimensional constellations
Without any additional structure that may reduce complexity, we must approximate for each $i=1,\ldots,d$, the conditional cdf $F({\theta_{i}}\!\mid\!\theta^{i-1})$ for each $\theta^{i-1}\in\mathcal{A}^{(i-1)}$ where $\mathcal{A}^{(i-1)}$ is the (i1)(i-1)-dimensional constellation previously constructed, whose size is $M^{i-1}$. This leads to a total complexity proportional to $1+M+\cdots+M^{d-1}=\Theta(M^{d})$. If, however, the model exhibits a Markovian structure of order $D$, then the approximation only needs to be performed for each $\theta_{i-D}^{i-1}$ in a $D$-dimensional grid with size $M^{D}$, yielding a reduced complexity of $\Theta(M^{D})$. In the fully general case, as the dimension $d$ grows, the complexity increases exponentially. Indeed, it is well known that sampling accurately from high dimension distributions is computationally hard in general.
@@ V-C Achievable rate for continuous output alphabet and approximate log-likelihood
$\mathbb{E}\Bigl[\nabla_{\boldsymbol{\theta}}\ln p_{\boldsymbol{\theta}}(Y\!\mid\!\hat{Y}=\hat{y})\,\nabla_{\boldsymbol{\theta}}\ln p_{\boldsymbol{\theta}}(Y\!\mid\!\hat{Y}=\hat{y})^{\top}\,\Big|\,\hat{Y}=\hat{y}\Bigr$]. (67)
@@ VII Conclusion
$C(P)$ $\geq I({\boldsymbol{\theta}};\boldsymbol{y})=\bar{C}(P)+o(1).$ (106)
@@
$C_{L}(P)-D(w_{J,\lambda^{*}}\|w_{J,\lambda_{L}^{*}})+\lambda_{L}^{*}\,\mathbb{E}_{w_{J,\lambda^{*}}}\!\big[c({\boldsymbol{\theta}})-P\big].$ (108)
@@
$\int_{\Theta}2^{-\lambda^{*}(c({\boldsymbol{\theta}})-P)}\sqrt{\det\boldsymbol{J}({\boldsymbol{\theta}})}\log\!\frac{\sqrt{\det\boldsymbol{J}({\boldsymbol{\theta}})}}{\sqrt{\det\boldsymbol{J}_{L}({\boldsymbol{\theta}})}}\,d{\boldsymbol{\theta}}\\ \leq\kappa_{0}\int_{\Theta}\ln\!\frac{\det\boldsymbol{J}({\boldsymbol{\theta}})}{\det\boldsymbol{J}_{L}({\boldsymbol{\theta}})}\,d{\boldsymbol{\theta}},$ (111)
with
$\kappa_{0}:=\sup_{{\boldsymbol{\theta}}\in\Theta}\left\{2^{-\lambda^{*}(c({\boldsymbol{\theta}})-P)}\sqrt{\det(\boldsymbol{J}({\boldsymbol{\theta}}))}\log e\right\}$ (112)
being a bounded constant since the function inside the supremum is continuous and the set $\Theta$ is compact. Plugging (107), (110), and (111) back in (108), we obtain an achievable rate $C(P)-\kappa_{0}e_{L}$ and complete the proof of the theorem.
@@
$=\nabla_{\boldsymbol{\theta}}\ln p_{\boldsymbol{\theta}}(y)-\frac{\int_{\mathcal{V}_{\ell}}\nabla_{\boldsymbol{\theta}}p_{\boldsymbol{\theta}}(u)\,du}{\int_{\mathcal{V}_{\ell}}p_{\boldsymbol{\theta}}(u)\,du}$ (118)
@@
$\boldsymbol{J}(Y;{\boldsymbol{\theta}}\!\mid\!\hat{Y}=\ell)\\ =\mathbb{E}\Bigl[\nabla_{\boldsymbol{\theta}}\ln p_{\boldsymbol{\theta}}(Y\!\mid\!\hat{Y}=\ell)\,\nabla_{\boldsymbol{\theta}}\ln p_{\boldsymbol{\theta}}(Y\!\mid\!\hat{Y}=\ell)^{\top}\,\Big|\,\hat{Y}$$\ell\Bigr$]. (123)
@@
$=36\rho^{2}\frac{\eta_{1}^{\prime}}{\eta_{1}\epsilon}\,L^{-\frac{2\eta_{1}\epsilon}{m\eta_{1}^{\prime}}}\left(\frac{\eta_{1}\epsilon\,K_{\epsilon}\kappa_{1}^{\epsilon/(2+\epsilon)}}{72(2+\epsilon)\rho^{2}}\right)^{\frac{2(2+\epsilon)}{\eta_{1}^{\prime}}},$ (134)
where $\eta_{1}^{\prime}:=\eta_{1}\epsilon+2(2+\epsilon)$. If $\sup_{{\boldsymbol{\theta}}\in\Theta}P_{{\boldsymbol{\theta}}}(\|Y\|>r)\leq\kappa_{2}e^{-\eta_{2}r^{2}}$ then